# Wyb. Wyspiańskiego 27, 50-370 Wrocław, Poland

### v1

@@ Abstract
Keywords:
problem complexity measures synthetic data genetic algorithm classification regression
@@ 1 Introduction
In today’s era of widespread applications of machine learning, considering both the risks and hopes associated with its rapid development and growing popularity [10], one must pay particular attention to the reliable and accurate evaluation of proposed methods. In many cases, using real-world datasets for method evaluation appears to confirm the effectiveness of a specific algorithm and demonstrates its usability [27]. However, a limited availability of diverse real-world datasets is a commonly recognized issue in machine learning research [16].
The use of synthetic data offers many benefits, including reduced storage requirements, since large amounts of data can be obtained based on a synthetic strategy and a particular random state. What is more, data with particular characteristics can be requested – i.e., with large dimensionality or specific relationships between features. For instance, a complex feature relationship can be found in the two moons dataset, commonly utilized for demonstration purposes in uncertainty quantification research [28].
When evaluating novel methodologies, showing an increase in the recognition ability for challenging datasets is usually seen as desirable. When using synthetic data, specifying hyperparameters of a generation method enables one to generate the data with various complexities. Such complexity may be related to the data dimensionality, the informativeness of features, the level of noise, or the class imbalance [7].
@@ 1.1 Synthetic datasets in machine learning research
One of the most frequently used classification data synthesis methodologies is Madelon [7], implemented in a make_classification function of scikit-learn library. The generator samples data distributions lying on the vertices of a multidimensional hypercube. A particular benefit of this approach is the possibility to specify the number of informative, redundant, and repeated features, as well as the length of the hypercube edge, related to the class separation. By modifying the hyperparameters of a generator, one can obtain datasets with diverse complexities.
Some other generation strategies focus on specific difficulties related to class overlaps across features. An example may be the already mentioned two moons dataset [28] or spiral and circles datasets, where classes are not linearly separable [1].
The synthetic data generation methods aim to produce datasets that verify certain capabilities of recognition methods. For example, Mitra et al. [22] proposed 10 different strategies for classification data generation to evaluate methods’ feature extraction abilities. Other research used data generated based on logic gates or LED displays [9]. Another reason for synthetic data usage is also to preserve confidential and sensitive information. In that case, some data generation methods aim to approximate the real-world sample distribution by employing smote or deep learning methods like Generative Adversarial Networks or Variational Autoencoders [5]. Such a generator aims to reliably resemble the actual real-world data [4].
In the case of a regression task, the synthetic data can be obtained by generating random features and then calculating the target variable according to their values. Friedman datasets [2] describe three types of non-linear relations between features and a target using trigonometric functions and feature multiplications. A particularly interesting synthetic dataset that can be used for regression and dimensionality reduction tasks is the swissroll dataset [20], which describes the points lying on a 3D nonlinear structure. A commonly used generation strategy, implemented in the make_regression function of the scikit-learn library, uses linear regressors on a random feature matrix and Gaussian noise with a specified level to produce the target variable.
As noted, the synthetic datasets aim to pose a challenge to the proposed algorithms, verifying their specific abilities. Scaling the difficulty of the generated data is a valuable possibility to reliably assess the proposed methods.
@@ 1.2 Complexity measures
Complexity measures are used to estimate the level of problem difficulty. Those measures have been proposed and systematized for classification [14] and regression tasks [15]. They usually serve as the tool of dataset description for experimental research [11], since various measures and their combinations can distinguish simple problems (e.g., linearly separable) from more complex ones [19, 26]. Some research uses complexity measures in the preprocessing stage. Such applications include the selection of data resampling ratios for imbalanced datasets [13] or noise filtering [25].
Following the benefits of synthetic data usage, the complexity metafeatures can serve as a valuable tool in the context of synthetic data generation. One of the first approaches of that kind used a genetic algorithm that searched for the combination of labels that optimized the fraction of borderline points (N1) complexity measure [17]. The research from 2018 used a greedy search algorithm to optimize F1, N1, and N3 measures individually [21], also by assigning existing samples to different classes. Later approaches focused on multi- and many-criteria optimization with sophisticated optimization algorithms like NSGA-III [3], extending the set of optimized measures to three (N1, N2, and F1) [18] or four (C1, L2, N1, F2) [6]. Some approaches adjusted the labels of real-world data distributions, aiming to achieve the specific complexity of a problem that resembled the complexity of real-world data [23]. In all of the mentioned research, the individuals in an evolutionary algorithm were the vectors describing the class labels of a specific sample distribution.
A different approach was proposed in Sy:Boid method [8], where authors adjusted the boids simulation, originally intended to model the movement of birds for computer graphics [24], to produce synthetic datasets with various complexities. This approach used a genetic algorithm to search through the hyperparameters of the generation method, so that the resulting dataset was characterized by a specific complexity. This approach optimized F1 and N1 measures.
@@ 1.3 Contribution
This work introduces a different approach to altering the complexity of a dataset, which modifies the existing data distribution with feature projections using a transformation matrix whose coefficients are optimized with the genetic algorithm. In the presented research, the source distribution is synthetically generated. By default, the dimensionality $d$ of a dataset remains unaffected, hence, each individual in the population is a matrix of size $d\times d$.
An example of feature projections using two different random transformation matrices $p_{0}$ and $p_{1}$ for a synthetic dataset with a dimensionality $d=10$ is presented in Figure 1. The first subplot presents the pca components of the original, unmodified dataset. The other two subplots present the principal components of a dataset after applying the random projection. The complexity measures calculated for the datasets, displayed in the figure titles, show that the complexity of the dataset increased after applying the projection.
Figure 1: The results of random feature projections to a 10-dimensional synthetic classification problem. The original data is shown in the first plot, and the results of feature transformation in the following ones.
Compared to existing strategies, the search space proposed in this approach is continuous and unbounded. An important implication of using feature transformations is the possibility to optimize regression tasks, since the algorithm itself does not affect the labels of an input dataset. The only modification in the case of regression complexity optimization is the selection of complexity measures dedicated to this particular task.
In summary, this work presents a genetic algorithm designed to transform datasets to a specified level of complexity. The proposed solution enables the use of a synthetically generated dataset for classification or regression tasks and allows adjusting the problem complexity level toward a desired target specified by hyperparameters. The main contributions of the presented research are as follows:
@@
  • Selection of complexity measures for considered tasks that show promising adjustment possibilities via feature projections.
  • Optimization of problem complexities of a synthetically generated datasets to five different levels using the proposed approach.
@@ 2 Method
This article presents a Evolutionary Projection-based Complexity Optimization (epco) – a genetic algorithm that optimizes a set of predefined criteria, $\mathcal{C}={C_{0},C_{1},C_{2},\dots,C_{N}}$ describing the complexity of a classification or regression problem, towards a target level specified with a hyperparameter. For each criterion $C_{i}$, a fitness function is defined as $F_{i}=|T_{i}-C_{i}|$, where $C_{i}$ is the value of the measure for a given individual, and $T_{i}$ is the desired target value of this measure. The goal is to minimize $F_{i}$.
@@ 2.1 Algorithm description
The operation of the proposed approach is presented in Algorithm 1. The method receives a source $d$-dimensional dataset ${X,y}$ as a parameter, as well as the set of complexity measures to optimize $\mathcal{C}$ with their target values $\mathcal{T}$. The remaining hyperparameters describe the evolutionary algorithm and express the population size $m$, number of iterations $I$, crossover and mutation ratios $r_{c}$ and $r_{m}$, as well as the decay factor $d$, which reduces the number of crossover and mutation operations at a later stage of the evolution.
@@
2:$\{X,y\}$ – source dataset, $\mathcal{C}$ – set of complexity measures to optimize, $\mathcal{T}$ – list of target values of measures, $m$ – population size, $I$ – number of iterations, $r_{c}$ – crossover ratio, $r_{m}$ – mutation ratio, $d_{f}$ – decay factor
@@
4:$\{X_{p},y\}$ – result dataset
5$\triangleright$ initialize population and measure fitness
6:$\mathcal{P}\leftarrow m$ matrices of size $d$ x $d$ with random values drawn from $\mathcal{N}(0,3)$
$\mathcal{F}\leftarrow\emptyset$
@@
14:    $\mathcal{F}\leftarrow$ order fitness according to $\mathcal{F}$triangleright crossover
@@
19:       $p^{\prime}\leftarrow crossover(s_{1},s_{2})$triangleright substitute individuals with worst fitness
@@
31:end $\triangleright$ return dataset transformed with leading projection
32:$X_{p}\leftarrow X\otimes p_{|\mathcal{C}|}$
@@
The procedure starts with initializing the population. Each individual is a transformation matrix of size $d\times d$, where $d$ describes the data dimensionality. The coefficients in transformation matrices are drawn from the normal distribution $\mathcal{N}(\mu,\sigma)$ with the mean value of $\mu=0$ and the standard deviation of $\sigma=3$. In the next stage of processing, the individuals are scored according to the absolute differences in criteria $\mathcal{C}$ and their target values $\mathcal{T}$. First, each projection in the population $p_{i}$ is used to produce a new feature set $X^{\prime}=X\otimes p_{i}$, where $\otimes$ describes matrix multiplication and $X$ the original features. Later, the fitness in a given criterion is calculated as $|T_{j}-C_{j}(X^{\prime},y)|$ for a measure $C_{j}$ and its target $T_{j}$. This is described in the lines 2:6 of the pseudocode.
@@
The optimization runs for a specified number of iterations. The size of a population remains constant, with the order of individuals in a population following their fitness score. The individuals are sorted according to fitness in each of the criteria $C_{i}$ and the sum of their fitness scores $\Sigma_{C}$. Then, the population is reordered in such a way that the first $|\mathcal{C}|$ individuals have the highest scores in consecutive criteria, and the individual in position $|\mathcal{C}|+1$ has the highest score in the sum of the criteria. The illustrative example of such an order is presented in Figure 2 for three complexity measures.
@@
Figure 2: The order of individuals in a population. The example presents the head of a population for three criteria: $C_{0}$, $C_{1}$, and $C_{2}$. Individuals are sorted according to fitness for each of the criteria and their sum $\Sigma_{C}$. Four leading individuals of the population have the highest fitness according to individual criteria and their sum.
Ordering individuals according to their fitness allows for a straightforward approach for selection by replacing the poorly adapted individuals in the tail of the population, and serves as a solution to select the best individuals for crossover by selecting the ones and the initial positions of the population.
@@
The crossover is described in the lines 10:17 of the pseudocode. The number of operations is calculated according to the crossover ratio $r_{c}$ and the size of the population $m$. At least one crossover will be performed in each iteration. The procedure selects one individual from the head of the population and another, randomly selected individual present in the population. The result of a crossover is a weighted sum of transformation matrices with random weights. By default, there is a higher weight associated with the better-adapted individual. The new individual is replacing the worst-adapted one from the tail of the population.
@@
Mutation is described with the lines 18:23 of the pseudocode. The operation is affecting a randomly selected individual in a population. During mutation, a transformation matrix is modified by adding a random value to each coefficient. The random values affecting the individual are drawn from a normal distribution $\mathcal{N}(0,0.1)$. The $\sigma=0.1$, which is significantly lower than the standard deviation of the distribution from which the individual’s values are originally drawn ($\sigma=3$), means that the mutation only slightly affects the transformation matrix.
After the crossover and mutation, the ratios $r_{c}$ and $r_{m}$ are reduced according to the decay factor $d_{f}$, which improves the stability of the search in the later evolution phase. With the significant decay factor and the number of iterations, the operation of the algorithm in the later phase of processing will involve just a single crossover operation per iteration.
@@
Figure 3: Classification measure optimization process. The example shows optimization of three complexity measures – F1, N1, and ClsCoef – with the first 3 subfigures showing the relation between each of the two criteria. The individual points indicate the fitness of individuals at a given optimization stage (depending on the color of a point – blue indicates initial iterations, red the final ones). $X$ markers indicate leaders of the population. The final plot shows the scores of leaders across 200 iterations.
@@
Figure 3 presents the optimization process of the epco algorithm for the exemplary set of three classification measures – F1, N1, and ClsCoef. The blue points indicate the score of individuals at the initial stage of the optimization, and the red points show the final population. The star marker shows the target of optimization – the point (0,0) indicating the achievement of a target value $T_{i}$ of a given complexity measure $C_{i}$. The last plot in the figure shows how the fitness of population leaders evolved throughout the evolution process in each of the criteria.
@@
The optimization loop continues until the iteration limit is reached. The result of an algorithm is a dataset ${X_{p},y}$ in which the problem features $X_{p}$ are a result of an original features’ projection with a specific individual from the population. By default, the individual with the highest score in all criteria ($\Sigma_{C}$) is selected for projection. If requested, the algorithm can return a dataset transformed by any projection from the population. For example, using the projection $p_{0}$ in the transformation in line 27 of the pseudocode will result in the selection of the individual with the highest fitness in the $C_{0}$ criterion.
@@
The proposition of a custom approach to optimization and not relying on NSGA-III [3], like known approaches that optimize more than three criteria [6], was motivated by enabling the strong diversification of results within the relatively small population and many criteria. In NSGA-III, the population is formed based on the reference points spanned across the pareto frontier. In epco, the optimization of individual criteria, as well as considering the sum of fitness among them, serves as an equivalent of $|\mathcal{C}|+1$ reference points. This is also a simple solution that is suitable for a large number of criteria (up to 10 in a classification task). In future research, the dataset generation approach relying on feature projections (characteristic of epco) can be extended to a more computationally costly, but more precise solution. This research aims to present a simple, yet effective solution to complexity optimization that relies on feature transformation, which is a novelty in comparison to the methods known from the literature.
@@
Table 1: The description of the complexity measures selected for optimization with the epco algorithm.
@@
S1 Output distribution – smoothness measure that calculates a difference in output values in a mst build on the data samples

### v2

@@ 1 Introduction
In today’s era of widespread applications of machine learning, considering both the risks and hopes associated with its rapid development and growing popularity [11], one must pay particular attention to the reliable and accurate evaluation of proposed methods. In many cases, using real-world datasets for method evaluation appears to confirm the effectiveness of a specific algorithm and demonstrates its usability [26]. However, the limited availability of diverse real-world datasets is a commonly recognized issue in machine learning research [17].
The use of synthetic data offers many benefits, including reduced storage requirements, as large amounts of data can be generated using a synthetic strategy and a specific random state. What is more, data with particular characteristics can be requested – i.e., with large dimensionality or specific relationships between features. When evaluating novel methodologies, demonstrating an increase in recognition performance on challenging datasets is usually considered desirable. Specifying hyperparameters of a generation method enables one to generate data that poses various challenges. Such challenges may be related to data dimensionality, feature informativeness, noise level, or class imbalance [8].
@@ 1.1 Synthetic datasets in machine learning research
One of the most frequently used classification data synthesis methodologies is Madelon [8], implemented in a make_classification function of scikit-learn library. The generator samples data distributions lying on the vertices of a multidimensional hypercube. A particular benefit of this approach is the possibility of specifying the number of informative, redundant, and repeated features, as well as the length of the hypercube edge related to class separation.
Some other generation strategies focus on specific difficulties related to class overlaps across features. An example may be the two moons dataset [27] or spiral and circles datasets, where classes are not linearly separable [1]. In general, synthetic data generation methods aim to produce datasets that test the capabilities of recognition methods. For example, Mitra et al. [22] proposed 10 strategies for generating classification data to evaluate methods’ feature-extraction abilities. Other research used data generated based on logic gates or LED displays [10]. Another reason for using synthetic data is to preserve confidential and sensitive information. In that case, some data generation methods aim to approximate the real-world sample distribution by employing smote or deep learning methods like Generative Adversarial Networks or Variational Autoencoders [6]. Such a generator aims to reliably resemble the actual real-world data [5].
In a regression task, synthetic data can be generated by randomly sampling features and then calculating the target variable based on their values. Friedman datasets [2] describe three types of non-linear relations between features and a target using trigonometric functions and feature multiplications. A particularly interesting synthetic dataset for regression and dimensionality reduction is the swissroll dataset [21], which describes points lying on a 3D non-linear structure. A commonly used generation strategy, implemented in the make_regression function of the scikit-learn library, uses linear regressors on a random feature matrix and Gaussian noise with a specified level to produce the target variable.
@@ 1.2 Complexity measures
Complexity measures aim to estimate the difficulty of a given recognition problem. Those measures have been proposed and systematized for classification [15] and regression tasks [16]. They usually serve as tools for dataset description in experimental research [12], since various measures and their combinations can distinguish simple problems (e.g., linearly separable) from more complex ones [20, 25]. Some research uses complexity measures in the preprocessing stage. Such applications include selecting data resampling ratios for imbalanced datasets [14] or noise filtering [24].
Given the benefits of synthetic data, the complexity metafeatures can serve as a valuable tool for data generation. One of the first approaches of that kind used a genetic algorithm that searched for the combination of labels that optimized the fraction of borderline points (N1) complexity measure [18]. The 2018 research used a greedy search algorithm to optimize the F1, N1, and N3 measures individually [3], also by assigning existing samples to different classes. Later approaches focused on multi- and many-criteria optimization with sophisticated optimization algorithms like NSGA-III [4], extending the set of optimized measures to three (N1, N2, and F1) [19] or four (C1, L2, N1, F2) [7]. Some approaches adjusted the labels of real-world data distributions to achieve the specific problem complexity that resembles that of real-world data [23]. In all of the mentioned research, the individuals in an evolutionary algorithm were the vectors describing the class labels of a specific sample distribution.
A different approach was proposed in the Sy:Boid method [9], where the authors adjusted the boids simulation, originally intended to model the movement of birds for computer graphics, to produce synthetic datasets with varying complexities. This approach used a genetic algorithm to search over the hyperparameters of the generation method, aiming to achieve a specified level of complexity. This approach optimized the F1 and N1 measures.
@@ 1.3 Contribution
This work introduces a different approach to altering the complexity of a dataset, which modifies the data distribution via feature projections using a transformation matrix whose coefficients are optimized via a genetic algorithm. In the presented research, the source distribution is synthetically generated. By default, the dataset’s dimensionality $d$ remains unchanged; hence, each individual in the population is a $d\times d$ transformation matrix.
Figure 1: Principal components of the dataset produced using random feature projections of a 10-dimensional synthetic classification problem. The original data is shown in the first plot, and the results of feature transformations in the following ones.
An example of feature projections using two different random transformation matrices $p_{0}$ and $p_{1}$ for a synthetic dataset with a dimensionality $d=10$ is presented in Figure 1. The first subplot presents the pca components of the original, unmodified dataset. The other two subplots show the principal components of a dataset after applying random projection. The complexity measures calculated for the datasets, displayed in the figure titles, indicate that the complexity of the dataset increased after applying the projection.
In summary, this work presents a genetic algorithm that transforms datasets to a specified level of complexity. The proposed solution enables the use of a synthetically generated dataset for classification or regression tasks and allows for adjusting the problem complexity level toward a desired target specified by hyperparameters. Compared to existing strategies, the search space proposed in this approach is continuous and unbounded.
The main contributions of the presented research are as follows:
@@
  • Selection of complexity measures for the considered tasks that show promising optimization possibilities via feature projections.
  • Transformation of synthetically generated datasets to five different complexity levels using the proposed approach.
@@ 2 Method
This article presents the Evolutionary Projection-based Complexity Optimization (epco) – a genetic algorithm that optimizes a set of predefined criteria, $\mathcal{C}=\{C_{0},C_{1},\dots,C_{N}\}$ describing the complexity of a classification or regression problem, towards a target level specified with a hyperparameter. For each criterion $C_{i}$, a fitness function is defined as $F_{i}=|T_{i}-C_{i}|$, where $C_{i}$ is the value of the measure for a given individual, and $T_{i}$ is the desired target value of this measure. The goal is to minimize $F_{i}$.
@@ 2.1 Algorithm description
The operation of the proposed approach is presented in Algorithm 1. The method receives a source $d$-dimensional dataset $\{X,y\}$ as a parameter, as well as the set of complexity measures to optimize $\mathcal{C}$ with their target values $\mathcal{T}$. The original dataset $X$ contains $n$ samples $X=\{x_{0},x_{1},\dots,x_{n-1}\}$, and each sample $x_{i}$ is a $d$-dimensional vector $x=[x^{0},x^{1},\dots,x^{d-1}]$. The $n$-dimensional label vector $y$ describes the continuous or discrete target for each sample $x_{i}$. The remaining hyperparameters describe the evolutionary algorithm: population size $m$, number of iterations $I$, crossover and mutation ratios $r_{c}$ and $r_{m}$, as well as the decay factor $d_{f}$, which reduces the number of crossover and mutation operations at a later stage of evolution.
@@
[…]
@@
[…]
@@
14:  $\mathcal{F}\leftarrow$ order fitness according to $\mathcal{F}$ $\triangleright$ crossover
@@
19:   $p^{\prime}\leftarrow crossover(s_{1},s_{2})$ $\triangleright$ substitute individuals with worst fitness
@@
[…]
@@
The procedure starts with initializing the population. Each individual is a transformation matrix of size $d\times d$, where $d$ describes the data dimensionality. The coefficients in transformation matrices are drawn from the normal distribution $\mathcal{N}(\mu,\sigma)$ with a mean value of $\mu=0$ and a standard deviation of $\sigma=3$. In the next stage of processing, the individuals are scored based on the absolute differences between the criteria $\mathcal{C}$ and their target values $\mathcal{T}$. First, each projection in the population $p_{i}$ is used to produce a new feature set $X^{\prime}=X\otimes p_{i}$, where $\otimes$ describes matrix multiplication and $X$ the original features. Later, the fitness in a given criterion is calculated as $|T_{j}-C_{j}(X^{\prime},y)|$ for a measure $C_{j}$ and its target $T_{j}$. This is described in the lines 2:6 of the pseudocode.
@@
The optimization runs for a specified number of iterations. The size of a population remains constant, with the order of individuals in a population following their fitness scores. The individuals are sorted according to fitness in each criterion $C_{i}$ and the sum of their fitness scores, $\Sigma_{C}$. Then, the population is reordered in such a way that the first $|\mathcal{C}|$ individuals have the highest scores in consecutive criteria, and the individual in position $|\mathcal{C}|+1$ has the highest score in the sum of the criteria. Ordering individuals by fitness allows a straightforward approach to selection, replacing poorly adapted individuals at the tail of the population, and serves as a solution for selecting the best individuals for crossover. The illustrative example of such an order is presented in Figure 2 for three complexity measures.
@@
Figure 2: The order of individuals in a population. The example presents the head of a population for three criteria: $C_{0}$, $C_{1}$, and $C_{2}$. Individuals are sorted according to fitness for each criterion and their sum, $\Sigma_{C}$. Four leading individuals of the population have the highest fitness according to individual criteria and their sum.
@@
The crossover is described in the lines 10:17 of the pseudocode. The number of operations is calculated based on the crossover ratio $r_{c}$ and the population size $m$. At least one crossover will be performed in each iteration. The procedure combines one individual from the head of the population with another, randomly chosen. The result of a crossover is a weighted sum of transformation matrices with random weights. By default, a higher weight is associated with the better-adapted individual. The new individual replaces the worst-adapted one from the tail of the population.
@@
The mutation is described with the lines 18:23 of the pseudocode. The operation affects a randomly selected individual. Mutation involves adding random noise to each coefficient. The random values affecting the transformation are drawn from a normal distribution $\mathcal{N}(0,0.1)$. The $\sigma=0.1$, which is significantly lower than the standard deviation of the distribution from which the individual’s values are originally drawn ($\sigma=3$), means that the mutation only slightly affects the transformation matrix. After the crossover and mutation, the ratios $r_{c}$ and $r_{m}$ are reduced according to the decay factor $d_{f}$, which improves the stability of the search in the later evolution phase. With the significant decay factor and the number of iterations, the algorithm’s operation in the later phases of processing will involve a single crossover per iteration.
@@
Figure 3: The optimization of three complexity measures – F1, N1, and ClsCoef. The first 3 subfigures show the relationships between the criterion pairs. The individual points indicate the population’s fitness at a given optimization stage. $X$ markers indicate leaders of the population. The last plot shows the leaders’ scores across 200 iterations.
@@
Figure 3 presents the optimization process of the epco algorithm for the exemplary set of three classification measures. The blue points indicate the scores of individuals at the initial stage of the optimization, while the red points show the final population. The star marker shows the target of optimization, indicating the achievement of a target value $T_{i}$ of a given complexity measure $C_{i}$. The last plot in the figure shows how the fitness of the population leaders evolved over the evolutionary process for each criterion.
@@
The optimization loop continues until the iteration limit is reached. The result of an algorithm is a dataset $\{X_{p},y\}$ in which the set of features $X_{p}$ results from the projection of original features’ with a specific individual from the population. By default, the individual with the highest score in all criteria ($\Sigma_{C}$) is selected for projection. If requested, the algorithm can return a dataset transformed by any projection from the population. For example, using the projection $p_{0}$ in the transformation in line 27 of the pseudocode will result in the selection of the individual with the highest fitness in the $C_{0}$ criterion.
@@
The proposal of a custom approach to optimization, rather than relying on NSGA-III [4, 7], was motivated by the potential for strong diversification within a relatively small population. In NSGA-III, the population is formed based on the reference points spanning the Pareto frontier. In contrast, epco optimized the individual criteria and their sum, serving as an equivalent of $|\mathcal{C}|+1$ reference points. In future research, the dataset generation approach relying on feature projections (characteristic of epco) can be extended to a more computationally costly but more precise algorithm. This research aims to present a simple yet effective solution to complexity optimization that relies on feature transformation, which is a novelty compared to methods known from the literature.
@@
Table 1: The description of the complexity measures selected for optimization with the epco algorithm
@@
S1 Output distribution – smoothness measure that calculates a difference in output values in a mst built on the data samples